Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 368     Entry time: 24 Sep 2023, 09:24
Author: Peter Zalden 
Group: FXE 
Subject: Shift summary 

7:00 Power glitch during SPB operation - they were supposed to hand the beam back at 7:30h

  • Power glitch brought down FXE interlock system with beam stop monitor alarm that could not be reset
  • SA1 PPL running fine, seems unaffected

9:10 Beam back in SA1, but still at 7 keV and cannot be changed by BKR to the req'd 5.8 keV, 94 kHz.

10:00 Beam back with FXE at req'd photon energy

Collected pump-probe diffraction and XAS data for the user expt. No further issues.

ELOG V3.1.4-7c3fd00